In order to obtain a 10-k you can do one of the following:

  1. visiting the SEC website at www.sec.gov and going to the EDGAR database (where you will look for the “10K”).
  2. on the company’s website
  3. calling the investor services or stockholder relations department of the company.

Name of companyTarget Corp

Principal exchange where the company trades _______________

Market price of the stock ________________ as of _____________ (date)

Annual dividend ________________________________________

Last dividend paid on ____________________________________

  1. Read the business summaries and management’s discussion and analysis (MD&A). Summarize the remarks (ex. Poor operating results/expansion plans/sale of a division, etc.)

  1. What kinds of products and/or services does the corporation provide/sell?
  1. a.         Which accounting/auditing firm audited the financial statements?
    1. Did the auditor’s report indicate any problems or exceptions?

5.         a.         Did you find any new terms within the financial statements?

b.         Did you find the presentation of the financial statements clear? Why or why not?

Part II

Indicate the formula to be used and insert the amounts used in computing each ratio. Express your answer in suitable units (percentages, decimals etc.)

  1. Current ratio
  1. Quick ratio
  1. Receivable turnover
  1. Average days sales uncollected
  1. Inventory turnover
  1. Profit margin
  1. Return on assets
  1. Return on Equity
  1. Earnings per share
  1. Debt to equity ratio
  1. Dividend yield
  1. Price earnings ratio (use price from question 1 and EPS from question 6-i)

Debt to Equity Ratio =

  Total Liabilities/Total Equity              100,000/200,000 = .5
Answer #1

Name of company Target Corp

Principal exchange where the company trades New York Stock Exchange

Market price of the stock $103 as of 28th February 2020 (date)

Annual dividend $1347 million (declared in 2018)

Last dividend paid on 2nd Feb 2018

Summarize the remarks (ex. Poor operating results/expansion plans/sale of a division, etc.): Steady increase in sales and earnings

What kinds of products and/or services does the corporation provide/sell: "everyday essentials and fashionable, differentiated merchandise at discounted prices."

Which accounting/auditing firm audited the financial statements: Ernst & Young LLP

Did the auditor’s report indicate any problems or exceptions: No

Did you find any new terms within the financial statements: No

Did you find the presentation of the financial statements clear? Why or why not? : Yes. The presentation was clear as it was in accordance with GAAP

Ratios: (amounts taken from from 10-k of Target Corp)

  1. Current ratio
    currentasset currentliability 12,519 15,014 = 0.833times
  1. Quick ratio
    currentasset - inventory currentliability 12,519 – 9,497 15,014 -=0.201times
  1. Receivable turnover
    NetCreditSales Average Accounts Receivables 75, 356 - 0 = N /A
  1. Average days sales uncollected
    0 Accounts Receivables Net Sales 75, 356 =N/A
  1. Inventory turnover

    Costof Goods Sold Average Inventory 53, 299 9047 = 5.891times
  1. Profit margin
    NetIncome Net Sales 2,937 75.356 = 4percent
  1. Return on assets
    Operating Income Total Assets 4, 110 41. 290
  1. Return on Equity
    NetIncome Shareholders Equity 2,937 11,297 00 = 26 percent
  1. Earnings per share
    NetIncome Weighted Average NumberOf Shares 2,937 518
  1. Debt to equity ratio
    Total Debt Total Equity 1,052 + 10, 223 -= 0.998times 11, 297
  1. Dividend yield
    DividendPer Share Market Price Per Share 2.54 103 a = 2percent
  1. Price earnings ratio
    Market PricePerShare Earnings Per Share 103 = 18.69 5.51